# Tricks from the Trade for Large-Scale Markdown Pricing: Heuristic Cut Generation for Lagrangian Decomposition

### v1

@@ 1 Introduction
Online retailers face challenges that are especially important in periods where economies stagnate and companies focus on efficiency: they need to guarantee a smooth experience throughout the whole customer journey, organize distribution of large assortments (in the millions of products) and efficiently manage inventory across multiple countries. A crucial aspect for the success of an online retailer is its capability to dynamically scale commercial processes to account for sales periods, traffic growth and external market conditions. This puts strict constraints on the underlying algorithms and systems, which need to deliver results in tight timelines (minutes or hours), even if suboptimal, to ensure business continuity. The large assortment necessitates a high degree of automation so that human oversight cannot be guaranteed for both scale and response time requirements. Key operational questions are hence fully automated. A prime example for such a system which is both operationally and commercially critical and is subject to scaling and response-time requirements is markdown pricing.
The goal in markdown pricing is to maximize profit while managing the fixed inventory for the entirety of a season.
In this paper, we consider large-scale price optimization systems for online retailers and the challenges that arise. Our show case throughout the paper will be Zalando SE, a leading online fashion retailer in Europe, throughout. With more than 50 million customers, presence in over 25 European countries and close to 2 million articles in its assortment, efficient pricing is one of the cornerstones of the Zalando business. To be successful, pricing systems and algorithms need to be
@@
In addition, markdown pricing is one of the most effective levers to boost various dimensions of commercial performance (discounts depth, revenue, profit, etc.), both in in specific countries/markets and on a company-wide level. This requires the pricing system to be able to steer towards strategic company goals, which in turn necessitates an integrated optimization (due to linking constraints) for all, or at least larger sets of markets. This poses an additional challenge for algorithm design, since the aforementioned commercial needs demand timely execution.
A typical approach in markdown pricing follows the predict-then-optimize paradigm (see e.g., (Ferreira et al., 2016; Loh et al., 2022; Kedia et al., 2020; Caro and Gallien, 2012) for examples in pricing) where a number of forecasting models (e.g., (Kunz et al., 2023; Schultz et al., 2023) for forecasting for pricing at Zalando and (Salinas et al., 2020; Lim et al., 2021; Rasul et al., 2021) for more general examples) provide input to an optimization problem. In practice, mixed integer programming (MIP) solutions are often employed given their adaptability to changing constraints (Kedia et al., 2020; Li et al., 2022; Caro and Gallien, 2012). The pricing approach of our show case Zalando follows this paradigm and here, we will consider in particular the ”optimize” part. Previous work (Li et al., 2022) described the basic setup of the pricing optimization problem at Zalando and key ingredients necessary for its solution, such as a Lagrangian decomposition framework, modelling approaches and basic heuristics. However, in the years since the deployment of this solution in production, roughly since 2021, there were occasional repercussions from unfavourable practical situations that led to non-convergence of the solution. As a further motivation for improving the heuristic performance of the markdown-pricing system, the scale of large online retailers like Zalando implies that even modest percentual gains in finding more profitable prices result in large absolute monetary gains – easily in the multi-digit million Euro range.
@@ 2 Related Work
Pricing and revenue management are mature areas of research stretching multiple disciplines, see e.g., (Phillips, 2021) for an introduction to the topic from an Economics perspective. The Machine Learning community typically covers aspects involving prediction, like demand forecasting, e.g., (Eisenach et al., 2020; Kunz et al., 2023; Oreshkin et al., 2020) and sometimes also taking a reinforcement learning perspective (e.g., (Madeka et al., 2022) for an example of an industrial application). Econometricians tend to focus on causal aspects or price elasticity estimation (Athey and Imbens, 2007; Strauss et al., 2018; de Chaisemartin and D’Haultfœuille, 2020; Deaton and Muellbauer, 1980; Fogarty, 2010; Hughes et al., 2008; DeFusco and Paciorek, 2017) and in the formulation of the optimization problems needed for profit-optimal pricing (Phillips, 2021). Here, we take an Operations Research erspective (see (Ferreira et al., 2016; Li et al., 2022) for examples), where we take both the predictive problems and the problem formulation as given and rather work on solving the optimization problem most efficiently within the requirements dictated by production purposes. This is, to the best of our knowledge, under-explored in this specific form in pricing. Chen et al. (2024) present a notable exception with a recent contribution, however, not for markdown pricing but for multi-product pricing in a ride-hauling setting.
@@ 3 Background and Problem Formulation
Additional business constraints restrict the solution space. Typical examples are targets for an average discount rate or a gross merchandising volume, both aggregated over each country. These targets are often refereed to as linking constraints, as they depend on all articles. In a general form, the primal pricing problem is thus given as
@@
where $Ax\leq b$ enforces the linking constraints and $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ describes the Cartesian product of the individual articles feasible sets.
@@
where $p_{i}$ denotes the un-discounted price of article $i$ and $s_{i,\bar{x}_{i}}$ the number of items sold of article $i$ when offered for price $\bar{x}_{i}$, and the set of variables $\bar{x}_{i}$ is a subset of the decision variables $x_{i}$.
@@ 4 Overview of Algorithms
The Cutting Plane approach avoids this problem by considering a relaxation to Problem (2), which contains only a subset of these potentially exponentially many constraints. The relaxation is tightened in every iteration by adding further valid inequalities. In more detail, assume that in iteration $j$ a number of inequalities determined by points $X^{1},\dots,X^{j}\in\mathcal{X}$ is included in the relaxation. Then the relaxation is given by
@@
is violated for $\mu^{j},\lambda^{j}$. The separation problem is equivalent to computing $\textnormal{LR}(\lambda^{j})$ and comparing its value against $\mu^{j}$. If $\textnormal{LR}(\lambda^{j})>\mu^{j}$, the maximizing solution of $\textnormal{LR}(\lambda^{j})$ is added as $X^{j+1}$. In the next iteration this leads to another inequality in the relaxation cutting off the current solution $\mu^{j},\lambda^{j}$. Otherwise the optimal Lagrangian solution value is found.
@@
To avoid that the relaxations given by (3) are unbounded during the first iterations, an upper bound $\lambda\leq\bar{\lambda}$ is added to the linear problem. This upper bound is also useful in case a linking constraint is infeasible. The corresponding multiplier would otherwise grow indefinitely. We choose the upper bound $\bar{\lambda}$ large enough that nearly all feasible linking constraints should have a non-violating solution to the Lagrangian relaxation with $\bar{\lambda}$. In consequence, we start the algorithm with $\lambda^{0}=0$, and it will choose $\lambda^{1}_{\ell}=\bar{\lambda}$ for all linking constraints $\ell$ violated by $X^{0}$. In cases with a large number of linking constraints (and consequently a high-dimensional Cutting Plane problem) it might require a large number of iterations and added cuts until all occurrences of $\lambda^{j}_{\ell}=\bar{\lambda}$ for at least one $\ell$ have been dropped from the cutting plane problem. Indeed, in practice we observe that most early iterations will contain multipliers at $\bar{\lambda}$, even though such multipliers over-fulfil on the constraints and are likely too high.
@@ 5 Heuristics for the Markdown Pricing Problem
In this section, we enhance the previously outlined general methodology by introducing a new strategy for identifying violated inequalities for the relaxation. As we will demonstrate, this enhancement not only improves the algorithm’s performance but also addresses potential convergence issues. This improvement can have a substantial impact on practical outcomes, leading to a significantly lower objective, which translates into improved long-term profitability.
The original Cutting Plane procedure obtains cuts by solving $\textnormal{LR}(\lambda)$. In our setting this involves heavy usage of (distributed) computational resources because determining objectives $f_{i}(x_{i})$ still is costly – each models the long term profit for one article for a season and includes the determination of prices and sales in various countries for all weeks of the season. Thus, we investigate heuristic methods to obtain valid cuts with less overhead, i.e., given $\lambda^{j}$ and $\mu^{j}$ we would like to find $x\in\mathcal{X}$ such that Inequality (4) is violated for the current iterations multipliers. Figure 1 illustrates this. The generation of such heuristic cuts has to trade-off speed vs. potential much stronger improvements by computing $\textnormal{LR}(\lambda^{j})$. Furthermore, changing the strategy has a down-stream impact as the evaluated solutions $X^{k}$ are used to construct the final solution, see the last part of the previous section.
The core of our idea is to efficiently combine past found solutions $X^{1},\dots,X^{j}$ for which $f_{i}(X^{k}_{i})$ is already computed for every iteration $k$ and article $i$ In that case we can find a new solution $X^{j+1}$ which is a combination of past solutions such that for all $i$ there is a $k\in{0,...,j}$ with $X_{i}^{j+1}=X^{k}_{i}$ Importantly, $X^{j+1}$ is a valid solution in $\mathcal{X}$, but not necessarily optimal for $\textnormal{LR}(\lambda,x)$ under any Lagrangian multiplier $\lambda$. Thus, there is no guarantee for finding a cut that maximizes the violation or even finding a violated cut at all.
@@
  • Random heuristic: For each article $i$ sample $X_{i}^{k}$ uniformly from $k\leq j$. This gives us a baseline to compare the other two strategies.
@@
  • Feasibility heuristic: Execute the primal heuristic from the previous section, using the current set of solutions (obtained from the preceding iterations). This process strives to derive a cut from a feasible and almost optimal solution. This approach is the most computationally involved among the three discussed strategies. Moreover, unlike the other two heuristics, repeatedly applying this strategy in sequence is not advantageous because it does not consider changes of the Lagrangian multipliers.
@@
We did not perform intense parameter tuning, the chosen values are mostly ad-hoc with some sensitivity tests. We limit the algorithm to $\bar{n}=10$ iterations. Up to $\bar{m}=10$ additional cuts could be generated. As tolerances, we ran our experiences with $\mbox{tol}_{e}=1$ for the efficacy and $\mbox{tol}_{\mu}=1\mathrm{e}{-6}$ for the dual bound gap.
@@ 6 Experiments and Real-world Impact
In the following section we illustrate the practical relevance of the methods described previously for the use case of the large-scale markdown price optimization system at Zalando. We report results here on a subset of the assortment for around 500,000 articles (depending on the exact time of the season), and for 14 countries, each with different discounts on weekly granularity until the season end of each article. Updated pricing recommendations are delivered at least twice a week, with 48 linking simplified discount rate (sDR) constraints submitted as targets (24 lower and upper bounds each). To integrate with Zalando’s business processes, the turn-around time between the setting of targets and the delivery of results should usually not exceed 3 hours.
Even though considerable work has been put into improving the speed at which we can solve the article problems for one set of Lagrangian multipliers since we last reported on it (Li et al., 2022), the time to solve the Lagrangian relaxation once remains at 150-700 CPU hours (using AWS C5 instances, exact time depending on the expected lifetime of the article). Even after parallelization over very large clusters the time to complete one iteration is therefore 3-12 minutes. Given additional overhead and because of both time and budgetary constraints we are limited to 10-15 iterations, a number that is far below the number of iterations one might usually expect for a problem with this number of constraints Guignard (2003).
In the experiments, we evaluate first which of the three cut generating strategies appears to be most effective, then the influence of the quality of the cuts on the dual bound. We demonstrate that the cut heuristics are not only faster but give a better trade-off for bound improvement over time. Finally, we show how the associated commercial key performance indicators (KPI) benefit from the approach presented here.
@@
To explore whether the heuristics described above help improve the quality of solutions, we first tested the relative effectiveness of the different strategies. Fig. 2 illustrates the performance via a case study of a problem with problematic convergence in the baseline (no-heuristic) strategy; other instances behave similarly. We find that the maximum violation heuristic is most effective, closing the relative dual gap, $\frac{\min_{k\leq j}\textnormal{LR}(\lambda^{k})-\mu^{j}}{\min_{k\leq j}\textnormal{LR}(\lambda^{k})}$, to less than 2% after six outer loop iterations and 0.4% in the 10th iteration for a problem which did not converge after 10 iterations in the baseline (with a relative gap much larger than 100%). We observed, however, that during the first few iterations, the relaxed primal bound hardly improved, and our imposed efficacy criteria would interrupt the heuristic cut generation early and continue with a full Lagrangian step.
Applying the feasibility heuristic is also effective in generating useful additional cuts, but less so. For that heuristic, the gap only drops below 100% in the 10th iteration, with a final gap of 7.4%. Random cuts are not effective, with the results being virtually indistinguishable from not applying any heuristic. We have observed similar results on other problem instances and concluded that the maximum violation heuristic strategy is performing the best. Thus we selected it as the most promising candidate for production and further results focus only on this heuristic.

### v2

@@ 1 Introduction
Online retailers face challenges that are especially important in periods where economies stagnate and companies focus on efficiency: they need to guarantee a smooth experience throughout the whole customer journey, organize distribution of large assortments (in the millions of products) and efficiently manage inventory across multiple countries. A crucial aspect for the success of an online retailer is its capability to dynamically scale commercial processes to account for sales periods, traffic growth and external market conditions. This puts strict constraints on the underlying algorithms and systems, which need to deliver results in tight timelines (minutes or hours), even if suboptimal, to ensure business continuity. The large assortment necessitates a high degree of automation so that human oversight cannot be guaranteed for both scale and response time requirements. Key operational questions are hence fully automated. A prime example for such a system that is both operationally and commercially critical and is subject to scaling and response-time requirements is markdown pricing. where The goal in markdown pricing is to maximize profit by managing a given inventory for the entire season.
In this paper, we consider large-scale price optimization systems for online retailers and the challenges that arise. Our show case throughout the paper will be Zalando SE, a leading online fashion retailer in Europe. With more than 50 million customers, presence in over 25 European countries and close to 2 million articles in its assortment, efficient pricing is one of the cornerstones of the Zalando business. To be successful, pricing systems and algorithms need to be
@@
In addition, markdown pricing is one of the most effective levers to boost various dimensions of commercial performance (discounts depth, revenue, profit, etc.), both in specific countries/markets and on a company-wide level. This requires the pricing system to be able to steer towards strategic company goals, which in turn necessitates an integrated optimization (due to linking constraints) for all, or at least larger sets of markets. This poses an additional challenge for algorithm design, since the aforementioned commercial needs demand timely execution.
A typical approach in markdown pricing follows the predict-then-optimize paradigm (see e.g., (Ferreira et al., 2016; Loh et al., 2022; Kedia et al., 2020; Caro and Gallien, 2012) for examples in pricing) where a number of forecasting models (e.g., (Kunz et al., 2023; Schultz et al., 2023) for forecasting for pricing at Zalando and (Salinas et al., 2020; Lim et al., 2021; Rasul et al., 2021) for more general examples) provide input to an optimization problem. In practice, mixed-integer programming (MIP) solutions are often employed given their adaptability to changing constraints (Kedia et al., 2020; Li et al., 2022; Caro and Gallien, 2012). The pricing approach of our show case Zalando follows this paradigm and here, we will consider in particular the ”optimize” part. Previous work (Li et al., 2022) described the basic setup of the pricing optimization problem at Zalando and key ingredients necessary for its solution, such as a Lagrangian decomposition framework, modelling approaches and basic heuristics. However, in the years since the deployment of this solution in production, roughly since 2021, there were occasional repercussions from unfavourable practical situations that led to non-convergence of the solution. As a further motivation for improving the heuristic performance of the markdown-pricing system, the scale of large online retailers like Zalando implies that even modest percentual gains in finding more profitable prices result in large absolute monetary gains – easily in the multi-digit million Euro range.
@@ 2 Related Work
Pricing and revenue management are mature areas of research stretching multiple disciplines, see e.g., (Phillips, 2021) for an introduction to the topic from an Economics perspective. The Machine Learning community typically covers aspects involving prediction, like demand forecasting, e.g., (Eisenach et al., 2020; Kunz et al., 2023; Oreshkin et al., 2020) and sometimes also taking a reinforcement learning perspective (e.g., (Madeka et al., 2022) for an example of an industrial application). Econometricians tend to focus on causal aspects or price elasticity estimation (Athey and Imbens, 2007; Strauss et al., 2018; de Chaisemartin and D’Haultfœuille, 2020; Deaton and Muellbauer, 1980; Fogarty, 2010; Hughes et al., 2008; DeFusco and Paciorek, 2017) and in the formulation of the optimization problems needed for profit-optimal pricing (Phillips, 2021). Here, we take an Operations Research perspective (see (Ferreira et al., 2016; Li et al., 2022) for examples), where we take both the predictive problems and the problem formulation as given and rather work on solving the optimization problem most efficiently within the requirements dictated by production purposes. This is, to the best of our knowledge, under-explored in this specific form in pricing. Chen et al. (2024) present a notable exception with a recent contribution, however, not for markdown pricing but for multi-product pricing in a ride-hauling setting.
@@ 3 Background and Problem Formulation
Additional business constraints restrict the solution space. Typical examples are targets for an average discount rate or a gross merchandising volume, both aggregated over each country. These targets are often referred to as linking constraints, as they depend on all articles. In a general form, the primal pricing problem is thus given as
@@
where $Ax\leq b$ enforces the linking constraints, the set $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ describes the Cartesian product of the individual articles feasible sets and $x=(x_{1},\dots,x_{n})$.
@@
where $p_{i}$ denotes the undiscounted price of article $i$ and $s_{i,\bar{x}_{i}}$ the number of items sold of article $i$ when offered for price $\bar{x}_{i}$, and the set of variables $\bar{x}_{i}$ is a subset of the decision variables $x_{i}$.
@@ 4 Overview of Algorithms
The Cutting Plane approach avoids this problem by considering a relaxation to Problem (2), which contains only a subset of these potentially exponentially many constraints. The relaxation is tightened in every iteration by adding further valid inequalities. In more detail, assume that in iteration $j$ a number of inequalities determined by points $X^{1},\dots,X^{j}\in\mathcal{X}$ are included in the relaxation. Then the relaxation is given by
@@
is violated for $\mu^{j},\lambda^{j}$. The separation problem is equivalent to computing $\textnormal{LR}(\lambda^{j})$ and comparing its value against $\mu^{j}$. If $\textnormal{LR}(\lambda^{j})>\mu^{j}$, the maximizing solution of $\textnormal{LR}(\lambda^{j})$ is added as $X^{j+1}$. In the next iteration this leads to another inequality in the relaxation cutting off the current solution $\mu^{j},\lambda^{j}$. Otherwise, the optimal Lagrangian solution value is found.
@@
To avoid that the relaxations given by (3) are unbounded during the first iterations, an upper bound $\lambda\leq\bar{\lambda}$ is added to the linear problem. This upper bound is also useful in case a linking constraint is infeasible. The corresponding multiplier would otherwise grow indefinitely. We choose the upper bound $\bar{\lambda}$ large enough that nearly all feasible linking constraints should have a non-violating solution to the Lagrangian relaxation with $\bar{\lambda}$. Consequently, we start the algorithm with $\lambda^{0}=0$, and it will choose $\lambda^{1}_{\ell}=\bar{\lambda}$ for all linking constraints $\ell$ violated by $X^{0}$. In cases with a large number of linking constraints (and consequently a high-dimensional Cutting Plane problem) it might require a large number of iterations and added cuts until all occurrences of $\lambda^{j}_{\ell}=\bar{\lambda}$ for at least one $\ell$ have been dropped from the cutting plane problem. Indeed, in practice, we observe that most early iterations will contain multipliers at $\bar{\lambda}$, even though such multipliers over-fulfill on the constraints and are likely too high.
@@ 5 Heuristics for the Markdown Pricing Problem
In this section, we enhance the previously outlined general methodology by introducing a new strategy for identifying violated inequalities for the relaxation. As we will demonstrate, this enhancement not only improves the algorithm’s performance but also addresses potential convergence issues. This improvement can have a substantial impact on practical outcomes, leading to a significantly better objective, which translates into improved long-term profitability.
The original Cutting Plane procedure obtains cuts by solving $\textnormal{LR}(\lambda)$. In our setting this involves heavy usage of (distributed) computational resources because determining objectives $f_{i}(x_{i})$ still is costly – each models the long-term profit for one article for a season and includes the determination of prices and sales in various countries for all weeks of the season. Thus, we investigate heuristic methods to obtain valid cuts with less overhead, i.e., given $\lambda^{j}$ and $\mu^{j}$ we would like to find $x\in\mathcal{X}$ such that Inequality (4) is violated for the current iterations multipliers. Figure 1 illustrates this. The generation of such heuristic cuts has to trade-off speed vs. potential much stronger improvements by computing $\textnormal{LR}(\lambda^{j})$. Furthermore, changing the strategy has a down-stream impact as the evaluated solutions $X^{k}$ are used to construct the final solution, see the last part of the previous section.
The core of our idea is to efficiently combine past found solutions $X^{1},\dots,X^{j}$ for which $f_{i}(X^{k}_{i})$ is already computed for every iteration $k$ and article $i$. In that case we can find a new solution $X^{j+1}$ which is a combination of past solutions such that for all $i$ there is a $k\in{0,\dots,j}$ with $X_{i}^{j+1}=X^{k}_{i}$. Importantly, $X^{j+1}$ is a valid solution in $\mathcal{X}$, but not necessarily optimal for $\textnormal{LR}(\lambda,x)$ under any Lagrangian multiplier $\lambda$. Thus, there is no guarantee for finding a cut that maximizes the violation or even finding a violated cut at all.
@@
  • Random heuristic: For each article $i$ sample $X_{i}^{k}$ uniformly from $k\leq j$. This gives us a baseline for comparing the other two strategies.
@@
  • Feasibility heuristic: Execute the primal heuristic from the previous section, using the current set of solutions (obtained from the preceding iterations). This process strives to derive a cut from a feasible and almost optimal solution. This approach is the most computationally involved among the three strategies discussed. Moreover, unlike the other two heuristics, repeatedly applying this strategy in sequence is not advantageous because it does not consider changes of the Lagrangian multipliers.
@@
We did not perform intense parameter tuning, the chosen values are mostly ad-hoc with some sensitivity tests. We limit the algorithm to $\bar{n}=10$ iterations. Up to $\bar{m}=100$ additional cuts could be generated. As tolerances, we ran our experiences with $\mbox{tol}_{e}=1$ for the efficacy and $\mbox{tol}_{\mu}=1\mathrm{e}{-6}$ for the dual bound gap.
Comparison with a disaggregated Cutting Plane formulation
The cut heuristics are based upon the special characteristic of the markdown pricing problem: It is decomposable, i.e., the feasible set is the Cartesian product of the individual article level feasible sets, $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$. This characteristic actually allows to also obtain a so-called disaggregated Guignard (2003); Frangioni (2005) or multi-cut Mitra et al. (2016) formulation
$\mu^{j},\lambda^{j},\nu^{j}=\operatorname*{arg\,min}_{\lambda\geq 0,\mu,\nu}$ $\mu+\lambda^{\top}b$ (8)
s.t. $\mu\geq\sum_{i\in N}\nu_{i},$
$\nu_{i}\geq f_{i}(X^{k}_{i})-\lambda^{\top}AX_{i}^{k}\quad\forall i\in N~{}\forall k\leq j.$
Compare this with the formulation to obtain a bound after applying the maximum violation cut heuristic $t$ times:
$\mu^{j},\lambda^{j}=\operatorname*{arg\,min}_{\lambda\geq 0,\mu}\mu+\lambda^{\top}b~{}\text{s.t.}~{}\mu$ $\geq\sum_{i\in N}f_{i}(X^{k}_{i})-\lambda^{\top}AX^{k},$ (9)
$\mu$ $\geq\sum_{i\in N}f_{i}(X^{k^{s}}_{i})-\lambda^{\top}AX_{i}^{k^{s}}~{}\forall 1\leq s<t,$
where $k^{s}$ indicates the maximizing indices in Formula 6 in round $s$ of applying the maximum violation heuristic.
One can show that the following strength guarantees on $\mu^{j}$ from the different optimization problems holds: The disaggregated formulation (8) is at least as strong as the aggregated formulation with heuristic cuts (9), which in turn is at least as strong as the aggregated formulation (3). The first follows as the heuristic cutting planes are implied from the extended formulation of (8) via the summation of individual constraints. The latter follows, since (3) contains fewer constraints.
However, the aggregated formulation (3) contains only $j+1$ variables and $j$ constraints compared to $j+s$ constraints with heuristic cuts and approximately $N+j$ variables and $Nj$ constraints of the disaggregated formulation (8). It turns out in our experiments, see Section (6), that solving (8) is not possible in practice due or very time consuming to its size. Thus using disaggregated cutting plane is not possible in our case. A compromise, which randomly partitions articles into $M$ groups $N=N_{1}\cup\dots\cup N_{M}$ to form the partially aggregated formulation
$\mu^{j},\lambda^{j},\nu^{j}=\operatorname*{arg\,min}_{\lambda\geq 0,\mu,\nu}$ $\mu+\lambda^{\top}b$ (10)
s.t. $\mu\geq\sum_{1\leq\ell\leq M}\nu_{\ell},$
$\nu_{\ell}\geq\sum_{i\in N_{\ell}}f_{i}(X^{k}_{i})-\lambda^{\top}AX_{i}^{k}\quad\forall 1\leq\ell\leq M~{}\forall k\leq j$
is therefore evaluated below.
One can furthermore show that after a finite (but worst-case $j^{N}$) number of cut heuristics application, the computed bound is equal to the bound from the disaggregated formulation. To see this, note that after projecting $\nu$ away in the disaggregated formulation, one obtains constraints of the form
$\mu\geq\sum_{i\in N}f_{i}(X^{k_{i}}_{i})-\lambda^{\top}AX_{i}^{k_{i}}\quad\forall i\in N~{}\forall k_{1},\dots,k_{N},k_{i}\leq j.$ (11)
Exactly this family of constraints is separated in the maximum violation cut heuristic.
@@ 6 Experiments and Real-world Impact
In the following section we illustrate the practical relevance of the methods described previously for the use case of the large-scale markdown price optimization system at Zalando. We report results here on a subset of the assortment for around 500,000 articles (depending on the exact time of the season), and for 14 countries, each with different discounts on weekly granularity until the season end of each article. Updated pricing recommendations are delivered at least twice a week, with 48 linking sales-weighted discount rate (sDR) constraints submitted as targets (24 lower and upper bounds each). To integrate with Zalando’s business processes, the turn-around time between the setting of targets and the delivery of results should usually not exceed 3 hours.
Although considerable work has been put into improving the speed at which we can solve the article problems for one set of Lagrangian multipliers since we last reported on it (Li et al., 2022), the time to solve the Lagrangian relaxation once remains at 150-700 CPU hours (using AWS C5 instances, exact time depending on the expected lifetime of the article). Even after parallelization over very large clusters the time to complete one iteration is therefore 3-12 minutes. Given additional overhead and because of both time and budgetary constraints we are limited to 10-15 iterations, a number that is far below the number of iterations one might usually expect for a problem with this number of constraints Guignard (2003).
In the experiments, we evaluate first which of the three cut generating strategies appears to be most effective, then the influence of the quality of the cuts on the dual bound. We go on to demonstrate that the cut heuristics are faster and easier to tune compared to partial aggregation strategies. Finally, we show how the associated commercial key performance indicators (KPI) benefit from the approach presented here.
@@
To explore whether the heuristics described above help improve the quality of solutions, we first tested the relative effectiveness of the different strategies. Fig. 2 illustrates the performance via a case study of a problem with problematic convergence in the baseline (no-heuristic) strategy; other instances behave similarly. We find that the maximum violation heuristic is most effective, closing the relative dual gap, $d_{j}=\frac{\min_{k\leq j}\textnormal{LR}(\lambda^{k})-\mu^{j}}{\min_{k\leq j}\textnormal{LR}(\lambda^{k})}$, to less than 2% after six outer loop iterations and 0.4% in the 10th iteration for a particularly hard problem which did not converge after 10 iterations in the baseline (with a relative gap much larger than 100%). We observed, however, that during the first few iterations, the relaxed primal bound hardly improved, and our imposed efficacy criteria would interrupt the heuristic cut generation early and continue with a full Lagrangian step.
Applying the feasibility heuristic is also effective in generating useful additional cuts, but less so. For that heuristic, the gap only drops below 100% in the 10th iteration, with a final gap of 7.4%. Random cuts are not effective, with the results being virtually indistinguishable from not applying any heuristic. We have observed similar results on other problem instances and concluded that the maximum violation heuristic strategy is performing the best. Thus we selected it as the most promising candidate for production and further results focus only on the maximum violation heuristic.